OpenStudy (anonymous):

is the point (1,-2) a solution to 4x-2y=8

jimthompson5910 (jim_thompson5910):

Let's find out

jimthompson5910 (jim_thompson5910):

(1,-2) means x = 1, y = -2

jimthompson5910 (jim_thompson5910):

4x-2y=8 4(1)-2y=8 ... replace x with 1, since x = 1 4(1) - 2(-2) = 8 ... replace y with -2, because y = -2 4 + 4 = 8 ... multiply 8 = 8 ... add The last equation is always true. So (1,-2) is indeed a solution to 4x-2y=8 because it satisfies the equation.

jimthompson5910 (jim_thompson5910):

Visually, the point (1,-2) lies on the graph of 4x-2y=8